OpenStudy (anonymous):

For these questions, state whether each system is consistent and independent, consistent and dependent, or inconsistent. 1. –x + y = –4 3x – 3y = 12 2. 2x – 5y = 8 15y – 6x = –24

OpenStudy (anonymous):

1. y=x-4 if you solve by substitution 2. y=5/2y+4 if you solve by substitution

OpenStudy (anonymous):

thats as far as I can do

hartnn (hartnn):

1. from 2nd equation, 3(x-y)=12 --> (x-y)=4 from 1st equation:-x+y=-4--->(x-y)=4 so basically both the equations are same and there will be no solution. no solution----> inconsistent system. clear with 1st ??

OpenStudy (anonymous):

Yes inconsistent for first

hartnn (hartnn):

oh wait i am sorry, there are infinitely many solutions for 1. infinitely many--->dependent and consistent.

OpenStudy (anonymous):

Ok cool, and whats the second one?

hartnn (hartnn):

2. from 2nd eq:15y-6x=-24 so, 5y-2x=-8 so 2x-5y=8 which is same as 1st equation. so this case is also same case as 1. so infinitely many solutions. so dependent and consistent .

OpenStudy (anonymous):

So consistent and dependent? You positive?

hartnn (hartnn):

yes,absolutely. did u get why infinite solutions?

OpenStudy (anonymous):

No, why?

hartnn (hartnn):

lets take 1. as example. both equations gave this :x-y=4 which means any values of x,y which satisfies x-y=4 are the solutions. like (5,1)(6,2)(7,3).....and so on till infinity. hence infinite solution,got it?